# JuggleRL: Mastering Ball Juggling with a Quadrotor via Deep Reinforcement Learning

### v1

@@ Abstract
Aerial robots interacting with objects must perform precise, contact-rich maneuvers under uncertainty. In this paper, we study the problem of aerial ball juggling using a quadrotor equipped with a racket, a task that demands accurate timing, stable control, and continuous adaptation. We propose JuggleRL, the first reinforcement learning-based system for aerial juggling. It learns closed-loop policies in large-scale simulation using systematic calibration of quadrotor and ball dynamics to reduce the sim-to-real gap. The training incorporates reward shaping to encourage racket-centered hits and sustained juggling, as well as domain randomization over ball position and coefficient of restitution to enhance robustness and transferability. The learned policy outputs mid-level commands executed by a low-level controller and is deployed zero-shot on real hardware, where an enhanced perception module with a lightweight communication protocol reduces delays in high-frequency state estimation and ensures real-time control. Experiments show that JuggleRL achieves an average of $311$ hits over $10$ consecutive trials in the real world, with a maximum of $462$ hits observed, far exceeding a model-based baseline that reaches at most $14$ hits with an average of $3.1$. Moreover, the policy generalizes to unseen conditions, successfully juggling a lighter 55 g ball with an average of $145.9$ hits. This work demonstrates that reinforcement learning can empower aerial robots with robust and stable control in dynamic interaction tasks.
@@ III-B Quadrotor Dynamics
A quadrotor is modelled as a 66 degree-of-freedom rigid body of mass $m$ and inertia matrix $\mathbf{I}=[\mathbf{I}_{x},\mathbf{I}_{y},\mathbf{I}_{z}]$. Its dynamics are modelled by:
@@ IV-C Reinforcement Learning with Generalization Enhancement
Action Space. The action space is 44-dimensional, representing the collective thrust and body rates (CTBR). The action at time step $t$ is defined as $\mathbf{a}_{t}=[c,\omega_{x},\omega_{y},\omega_{z}]\in\mathbb{R}^{4}$, where $c$ denotes the collective thrust controlling the quadrotor’s vertical acceleration, and $(\omega_{x},\omega_{y},\omega_{z})$ are the body rates around the roll, pitch, and yaw axes, respectively.
@@ V-A Experiment Setting
We employ OmniDrones [15], a high-parallel UAV simulator, to construct the ball juggling task. The setup consists of a quadrotor (mass $1090$ g) juggling a ball (mass $47.2$ g, radius 22 cm) falling from above, using a racket with a $5.5$ cm radius. Each episode is capped at $500$ steps. Training is conducted with $4096$ parallel environments running at $50$ Hz, producing 22 billion simulation frames in under 33 hours on a desktop workstation (Intel i7-13700K CPU, RTX $4090$ GPU).
@@ V-D Main Results
Simulation Results. We first evaluate the performance of JuggleRL against MBPP in simulation. Fig. 4 compares different initial ball release heights. MBPP requires a fixed target apex height (set to 33 m in our experiments) in order to compute a feasible trajectory, whereas JuggleRL operates within a target apex height range of $[2.85,3.2]$ m without such constraints. MBPP succeeds only when the ball is released between $4.0$ and $4.5$ m, and fails completely at lower heights, primarily because it relies on sufficient reaction time for the prediction–planning pipeline. In contrast, JuggleRL consistently maintains near-maximum episode performance (over $58$ hits) across all tested heights, demonstrating its reactive capability and robustness.
@@ V-F Generalization
We further evaluate zero-shot transfer to an out-of-distribution setting: juggling a lighter 55g ball. As shown in Tab.˜VI, the policy achieves an average of $145.9$ hits over $10$ consecutive trials, with results ranging from 1919 to $494$. The high variance arises from the amplified sensitivity of the lighter ball to small control errors, yet the upper bound demonstrates that the learned strategy generalizes beyond the training distribution.
@@ VI Conclusion and Future Work
We propose JuggleRL, a quadrotor-based ball juggling system designed for dynamic physical interaction. To enable robust sim-to-real transfer, the system integrates (i) high-fidelity simulation through system identification, (ii) carefully designed reward shaping and domain randomization, and (iii) a lightweight communication protocol that mitigates observation latency. With this design, JuggleRL achieves up to $462$ consecutive hits in real-world experiments, dramatically outperforming a model-based baseline that manages at most $14$ hits. In addition, the system demonstrates out-of-distribution generalization, successfully juggling a lighter 55 g ball with an average of $145.9$ hits, despite its higher sensitivity to control errors.

### v2

@@ Abstract
Aerial robots interacting with objects must perform precise, contact-rich maneuvers under uncertainty. In this paper, we study the problem of aerial ball juggling using a quadrotor equipped with a racket, a task that demands accurate timing, stable control, and continuous adaptation. We propose JuggleRL, the first reinforcement learning-based system for aerial juggling. It learns closed-loop policies in large-scale simulation using systematic calibration of quadrotor and ball dynamics to reduce the sim-to-real gap. The training incorporates reward shaping to encourage racket-centered hits and sustained juggling, as well as domain randomization over ball position and coefficient of restitution to enhance robustness and transferability. The learned policy outputs mid-level commands executed by a low-level controller and is deployed zero-shot on real hardware, where an enhanced perception module with a lightweight communication protocol reduces delays in high-frequency state estimation and ensures real-time control. Experiments show that JuggleRL achieves an average of $311$ hits over $10$ consecutive trials in the real world, with a maximum of $462$ hits observed, far exceeding a model-based baseline that reaches at most $14$ hits with an average of $3.1$. Moreover, the policy generalizes to unseen conditions, successfully juggling a lighter $5$ g ball with an average of $145.9$ hits. This work demonstrates that reinforcement learning can empower aerial robots with robust and stable control in dynamic interaction tasks.
@@ III-B Quadrotor Dynamics
A quadrotor is modelled as a $6$ degree-of-freedom rigid body of mass $m$ and inertia matrix $\mathbf{I}=[\mathbf{I}_{x},\mathbf{I}_{y},\mathbf{I}_{z}]$. Its dynamics are modelled by:
@@ IV-C Reinforcement Learning with Generalization Enhancement
Action Space. The action space is $4$-dimensional, representing the collective thrust and body rates (CTBR). The action at time step $t$ is defined as $\mathbf{a}_{t}=[c,\omega_{x},\omega_{y},\omega_{z}]\in\mathbb{R}^{4}$, where $c$ denotes the collective thrust controlling the quadrotor’s vertical acceleration, and $(\omega_{x},\omega_{y},\omega_{z})$ are the body rates around the roll, pitch, and yaw axes, respectively.
@@ V-A Experiment Setting
We employ OmniDrones [15], a high-parallel UAV simulator, to construct the ball juggling task. The setup consists of a quadrotor (mass $1090$ g) juggling a ball (mass $47.2$ g, radius $2$ cm) falling from above, using a racket with a $5.5$ cm radius. Each episode is capped at $500$ steps. Training is conducted with $4096$ parallel environments running at $50$ Hz, producing $2$ billion simulation frames in under $3$ hours on a desktop workstation (Intel i7-13700K CPU, RTX $4090$ GPU).
@@ V-D Main Results
Simulation Results. We first evaluate the performance of JuggleRL against MBPP in simulation. Fig. 4 compares different initial ball release heights. MBPP requires a fixed target apex height (set to $3$ m in our experiments) in order to compute a feasible trajectory, whereas JuggleRL operates within a target apex height range of $[2.85,3.2]$ m without such constraints. MBPP succeeds only when the ball is released between $4.0$ and $4.5$ m, and fails completely at lower heights, primarily because it relies on sufficient reaction time for the prediction–planning pipeline. In contrast, JuggleRL consistently maintains near-maximum episode performance (over $58$ hits) across all tested heights, demonstrating its reactive capability and robustness.
@@ V-F Generalization
We further evaluate zero-shot transfer to an out-of-distribution setting: juggling a lighter 55g ball. As shown in Tab.˜VI, the policy achieves an average of $145.9$ hits over $10$ consecutive trials, with results ranging from $19$ to $494$. The high variance arises from the amplified sensitivity of the lighter ball to small control errors, yet the upper bound demonstrates that the learned strategy generalizes beyond the training distribution.
@@ VI Conclusion and Future Work
We propose JuggleRL, a quadrotor-based ball juggling system designed for dynamic physical interaction. To enable robust sim-to-real transfer, the system integrates (i) high-fidelity simulation through system identification, (ii) carefully designed reward shaping and domain randomization, and (iii) a lightweight communication protocol that mitigates observation latency. With this design, JuggleRL achieves up to $462$ consecutive hits in real-world experiments, dramatically outperforming a model-based baseline that manages at most $14$ hits. In addition, the system demonstrates out-of-distribution generalization, successfully juggling a lighter $5$ g ball with an average of $145.9$ hits, despite its higher sensitivity to control errors.